# Mr. DETR: Instructive Multi-Route Training for Detection Transformers

### v3

@@ 3.2 Multi-route Training
Auxiliary Route with Instructive Self-Attention. In Sec. 1, we discuss the motivation for incorporating independent self-attention in Route-3 as shown in Fig. 3(e), designed for one-to-many prediction. To further reduce trainable parameters and enhance parameter sharing with the primary route, we propose an innovative instruction mechanism as shown in Fig. 4. This mechanism guides shared object queries to enable one-to-many prediction. The query output $\hat{\mathbf{Q}_{3}}$ of Route-3 is written as:
@@
3.3 Instructive Self-Attention
As described in Sec. 3.2, Route-3 is realized using an instructive self-attention, which incorporates learnable instruction tokens $\mathbf{Q^{ins}}$ into object queries $\mathbf{Q}$, creating a combined sequence $\mathbf{\hat{\mathbf{Q}}^{ins}}$. Self-Attention is performed on this combined sequence. In this section, we examine distinct implementations of instructive self-attention aimed at informing object queries to realize one-to-many predictions, sharing self-attention parameters with the primary route. The simplest approach, depicted in Fig. 5(a), involves using a separate set of queries as inputs to facilitate one-to-many predictions. To improve the compatibility of object queries across different routes, we introduce learnable tokens that serve as instruction, delineating the prediction objectives. As depicted in Fig. 5(b), these instruction tokens are incorporated into shared object queries through addition. This approach necessitates a fixed number of instruction tokens equivalent to the query count. Unlike the addition approach, our method adapts instruction tokens through concatenation, thereby providing greater flexibility. As illustrated in Fig. 5(c), this flexibility extends not just to the number of instruction tokens, but also allows these learned tokens to dynamically transmit information to object queries via self-attention. Self-attention is performed on the combined sequence $\mathbf{\hat{\mathbf{Q}}^{ins}}$, but the corresponding outputs of the instructive tokens are discarded after self-attention because they are not intended for object localization. Further details can be found in Eqn. (6). In this way, the instruction tokens provide effective and adaptable guidance to Route-3, enabling it to make one-to-many predictions while utilizing shared parameters with the primary route. In Sec. 4.5, we evaluate the performance of various implementations.
@@ 3.3 Instructive Self-Attention
Specifically, we build $m$ learnable tokens $\mathbf{Q^{ins}}=\{q_{0}^{ins},q_{1}^{ins},...,q_{m-1}^{ins}\}$, called instruction tokens. Initially, these instruction tokens are attached to the input sequence of self-attention by concatenation, forming a composite set of input queries $\mathbf{\hat{Q}^{ins}}=\{q_{0},q_{1},...,q_{n-1},q_{0}^{ins},q_{1}^{ins},...,q_{m-1}^{ins}\}$, resulting in a length of $n+m$. Subsequently, self-attention is performed on these combined queries. As the attached instruction tokens are not utilized for object localization, their outputs are discarded post self-attention, serving solely to convey information to object queries. The output $\hat{\mathbf{Q}}_{3}$ of Route-3 is written as:
$\hat{\mathbf{Q}_{3}}=(\texttt{FFN}_{o2o}\circ\texttt{CA}\circ\texttt{E}\circ\texttt{SA})(\mathbf{\hat{\mathbf{Q}}^{ins}}),$ (6)
@@
where the function E eliminates the outputs of instruction tokens after self-attention. Notably, all components of Route-3, including self-attention, cross-attention, and FFN, share parameters with the primary route. Instruction tokens effectively guide object queries and subsequent modules to achieve one-to-many prediction. The shared parameters between the auxiliary and primary routes benefit the one-to-one prediction of the primary route.
@@
Discussion.  Large language models have demonstrated impressive capabilities [40] when equipped with meticulously crafted prompts [49, 23]. Subsequent researches treat these prompts as learnable tokens, achieving success in both natural language processing [29, 26, 36] and computer vision applications [69, 24, 54]. While prompt tuning [22, 48, 58] and our proposed instructive self-attention both involve attaching learnable tokens to the input sequence, our instruction token uniquely facilitates distinguishing between one-to-one and one-to-many assignments, rather than merely fine-tuning the model. Additionally, different from the typical Mixture-of-Experts methods [47, 11] that dynamically select experts based on the input feature through a gating mechanism, our separated FFN pre-assigns an expert to each specific task.
Model w/ Mr. DETR Epochs Queries mAP AP50 AP75 APs APm APl
@@
Table 2: The performance on the COCO 2017 [32] validation set. All models are based on the ResNet-50 [17] backbone.
4 Experiments
4.1 Setup
@@
Dataset and Evaluation. In line with existing research [70, 61, 21], we conduct experiments on the COCO 2017 [32], providing results in terms of standardized metrics, namely average precision (mAP, AP50, AP75) at various IoU thresholds. APs, APm, and APl refer to the average precision for small, medium, and large objects, respectively. For all experiments with 300 and 900 queries, we report performance according to top-100 and top-300 predictions, respectively.
Model Epochs Queries mAP AP50 AP75 APs APm APl
H-DETR [21] 12 300 55.9 75.2 61.0 39.1 59.9 72.2
DINO [61] 12 900 56.8 75.4 62.3 41.1 60.6 73.5
Co-DETR [71] 12 900 56.9 75.5 62.6 40.1 61.2 73.3
Salience-DETR [18] 12 900 56.5 75.0 61.5 40.2 61.2 72.8
Rank-DINO [43] 12 900 57.6 76.0 63.4 41.6 61.4 73.8
DAC-DINO [20] 12 900 57.3 75.7 62.7 40.1 61.5 74.4
EASE-DINO [14] 12 900 57.8 76.7 63.3 40.7 61.9 73.7
Stable-DINO [35] 12 900 57.7 75.7 63.4 39.8 62.0 74.7
Relation-DINO [19] 12 900 57.8 76.1 62.9 41.2 62.1 74.4
Mr. DETR (ours) 12 900 58.4 76.3 63.9 40.8 62.8 75.3
Table 3: The performance on the COCO 2017 [32] validation set. All models use the Swin-L [38] as the architecture backbone. In this experiment, Mr. DETR is built upon the DINO detection framework [61] with IA-BCE loss [1]. The denotes the results derived from [35].
@@
Implement Details. We conduct all experiments based on the ResNet-50 [17] and Swin-L [38] as the backbone pretrained on the ImageNet [10]. In all experiments, we apply the same data augmentations as [70, 61, 21, 20, 65]. AdamW [39] is employed as the optimizer, with the initial learning rate and weight decay established at 2e-4 and 1e-4, respectively. We utilize a batch size of 16 for training all the models. During training schedules of 12 and 24 epochs, the learning rate is reduced by a factor of 0.1 following the 11th and 20th epochs, respectively. Within the one-to-many assigner, the hyper-parameters are configured as follows: $K=6$, $\alpha=0.3$, and $\tau=0.4$. In all experiments with instructive self-attention, we set the number of instruction tokens as 10. For models based on DINO [61], we use 100 contrastive denoising queries.
4.2 Main Results
We conduct extensive experiments on the COCO 2017 dataset [32]. As shown in Tab. 2, in the training schedule of 12 epochs, our method achieves consistent improvement across various baseline models. Extending to 24 epochs, our approach enables Deformable-DETR++ [70] to achieve a 51.4% mAP, outperforming the baseline by 1.6%. Compared to other variants with one-to-many auxiliary training, our method exceeds DAC-DETR [20] and MS-DETR [65] by 0.9% and 0.5%, respectively. Using our approach, DINO baseline [61] can reach 50.9% mAP over 12 epochs, outperforming variants like Group-DETR [5], DAC-DETR [20] and MS-DETR [65]. To demonstrate our method’s compatibility with localization-aware loss [60, 1], we apply it to Align-DETR [1], achieving gains of 1.2% and 1.0% mAP in 12 and 24 epoch schedules, respectively. Our experiments consistently demonstrate improvements across different baseline models. Further experiments with the Swin-L [38] backbone, as shown in Tab. 3, reveal that our method achieves a 58.4% mAP in a 12-epoch training schedule, surpassing other variants. These experimental results underscore the effectiveness of our proposed method.
Epochs w/ Mr. DETR Mask mAP Box mAP
12 32.4 46.5
12 36.0 (+3.6) 49.5 (+3.0)
24 35.1 48.6
24 37.6 (+2.5) 50.3 (+1.7)
@@ 4.2 Main Results
Table 4: Instance segmentation results on the COCO 2017 [32] validation set. All experiments are based on Deformable-DETR++ [70] with 300 queries and ResNet-50 [17] as backbone.
4.3 Extension to Instance Segmentation

### v4

@@ III-B Multi-route Training
Auxiliary Route with Instructive Self-Attention. In Sec. I, we discuss the motivation for incorporating independent self-attention in Route-3 as shown in Fig. 3(e), designed for one-to-many prediction. To further reduce trainable parameters and enhance parameter sharing with the primary route, we propose an innovative instruction mechanism as shown in Fig. 4. This mechanism guides shared object queries to enable one-to-many prediction. The query output $\hat{\mathbf{Q}_{3}}$ of Route-3 is written as:
@@
III-C Instructive Self-Attention
As described in Sec. III-B, Route-3 is realized using an instructive self-attention, which incorporates learnable instruction tokens $\mathbf{Q^{ins}}$ into object queries $\mathbf{Q}$, creating a combined sequence $\mathbf{\hat{\mathbf{Q}}^{ins}}$. Self-Attention is performed on this combined sequence. In this section, we examine distinct implementations of instructive self-attention aimed at informing object queries to realize one-to-many predictions, sharing self-attention parameters with the primary route. The simplest approach, depicted in Fig. 5(a), involves using a separate set of queries as inputs to facilitate one-to-many predictions. To improve the compatibility of object queries across different routes, we introduce learnable tokens that serve as instruction, delineating the prediction objectives. As depicted in Fig. 5(b), these instruction tokens are incorporated into shared object queries through addition. This approach necessitates a fixed number of instruction tokens equivalent to the query count. Unlike the addition approach, our method adapts instruction tokens through concatenation, thereby providing greater flexibility. As illustrated in Fig. 5(c), this flexibility extends not just to the number of instruction tokens, but also allows these learned tokens to dynamically transmit information to object queries via self-attention. Self-attention is performed on the combined sequence $\mathbf{\hat{\mathbf{Q}}^{ins}}$, but the corresponding outputs of the instructive tokens are discarded after self-attention because they are not intended for object localization. Further details can be found in Eqn. (6). In this way, the instruction tokens provide effective and adaptable guidance to Route-3, enabling it to make one-to-many predictions while utilizing shared parameters with the primary route. In Sec. IV-F, we evaluate the performance of various implementations.
@@ III-C Instructive Self-Attention
Specifically, we build $m$ learnable tokens $\mathbf{Q^{ins}}=\{q_{0}^{ins},q_{1}^{ins},...,q_{m-1}^{ins}\}$, called instruction tokens. Initially, these instruction tokens are attached to the input sequence of self-attention by concatenation, forming a composite set of input queries $\mathbf{\hat{Q}^{ins}}=\{q_{0}^{ins},q_{1}^{ins},...,q_{m-1}^{ins},q_{0},q_{1},...,q_{n-1}\}$, resulting in a length of $m+n$. Subsequently, self-attention is performed on these combined queries. As the attached instruction tokens are not utilized for object localization, their outputs are discarded post self-attention, serving solely to convey information to object queries. The output $\hat{\mathbf{Q}}_{3}$ of Route-3 is written as:
$\hat{\mathbf{Q}_{3}}=(\texttt{FFN}_{o2o}\circ\texttt{CA}\circ\texttt{R}\circ\texttt{SA})(\mathbf{\hat{\mathbf{Q}}^{ins}}),$ (6)
@@
where the function R eliminates the outputs of instruction tokens after self-attention. Notably, all components of Route-3, including self-attention, cross-attention, and FFN, share parameters with the primary route. Instruction tokens effectively guide object queries and subsequent modules to achieve one-to-many prediction. The shared parameters between the auxiliary and primary routes benefit the one-to-one prediction of the primary route.
@@
Discussion.  Large language models have demonstrated impressive capabilities [77] when equipped with meticulously crafted prompts [78, 79]. Subsequent researches treat these prompts as learnable tokens, achieving success in both natural language processing [80, 81, 82] and computer vision applications [83, 84, 85, 86]. While prompt tuning [87, 88, 89] and our proposed instructive self-attention both involve attaching learnable tokens to the input sequence, our instruction token uniquely facilitates distinguishing between one-to-one and one-to-many assignments, rather than merely fine-tuning the model. We name our method Mr. DETR, due to its Multi-route training nature for DETR-based detectors.
Figure 6: Illustration of the proposed Mr. DETR++ architecture. Our approach replaces the two independent FFNs in Mr. DETR with a route-aware MoE, enabling knowledge sharing and task-specific specialization across routes. Two distinct gating functions, G and G’, are employed: G governs Route-1, while G’ is shared between Route-2 and Route-3. The MoE block is further integrated into the transformer encoder. To reduce computational cost, the MoE block is applied to low-scale features, with all tokens sharing a single expert.
III-D From Mr. DETR to Mr. DETR++
As illustrated in Fig. 3(g), to enable knowledge sharing between the two independent FFNs used in Mr. DETR, we propose a route-aware MoE to replace the FFNs. We further extend the MoE architecture to the transformer encoder to enhance feature representation. To mitigate computational overhead, we introduce a scale-aware MoE applied to low-scale features. Additionally, we incorporate a localization-aware score to improve classification calibration. This extended method is called Mr. DETR++.
[…]
Formally, the query output $\hat{\mathbf{Q}_{1}}$ for Route-1 in Mr. DETR++ is expressed as:
$\hat{\mathbf{Q}_{1}}=(\texttt{E}\circ\texttt{G'}\circ\texttt{CA}\circ\texttt{SA})(\mathbf{Q}),$ (7)
where SA denotes self-attention, CA denotes cross-attention, G’ is the gating function, and E represents the expert computation. Similarly, the query output $\hat{\mathbf{Q}_{2}}$ for the one-to-one prediction in Route-2 is given by:
$\hat{\mathbf{Q}_{2}}=(\texttt{E}\circ\texttt{G}\circ\texttt{CA}\circ\texttt{SA})(\mathbf{Q}).$ (8)
For Route-3, equipped with instructive self-attention, the query output $\hat{\mathbf{Q}_{3}}$ is:
$\hat{\mathbf{Q}_{3}}=(\texttt{E}\circ\texttt{G}\circ\texttt{CA}\circ\texttt{InstructSA})(\mathbf{Q}).$ (9)
This route-aware MoE design facilitates knowledge sharing while mitigating potential conflicts.
[…]
$\mathcal{T}_{k}=\text{TopK}(\mathbf{s},k),$ (10)
where $\text{TopK}(\mathbf{s},k)$ returns the indices of the $k$ experts with the highest scores. The output of the MoE block for the query $\mathbf{Q}$ is then computed as a weighted sum of the selected experts’ outputs:
$\hat{\mathbf{Q}}=\sum_{i\in\mathcal{T}_{k}}c_{i}\cdot e_{i}(\mathbf{Q}),$ (11)
where $e_{i}(\mathbf{Q})$ is the output of expert $e_{i}$ for the input query $\mathbf{Q}$, and $c_{i}$ is the corresponding gating score.
Scale-aware Mixture-of-Experts in Encoder. We further extend the MoE framework to the transformer encoder to enhance feature representations. Typically, the encoder processes a sequence of flattened feature tokens from multiple scales, $F=\{f_{0},f_{1},\dots,f_{d-1}\}$, where $f_{i}$ represents image tokens at scale $i$. Directly applying MoE to this long sequence would significantly increase computational cost. To address this, we propose a scale-aware MoE that applies a shared expert to all tokens and a specialized MoE block to low-scale features.
Assuming the token sequence length increases with scale index, similar to Eqn. (11), we apply the MoE to the first $\eta$ scales (low-scale features), and a shared FFN is applied to all tokens in $F$. The final features are computed by combining outputs for the first $\eta$ scales and retaining the FFN output for higher scales:
$\hat{f}_{j}^{\prime}=\begin{cases}\text{FFN}(f_{j})+\text{MoE}(f_{j}),&j=0,1,\dots,\eta-1,\\ \text{FFN}(f_{j}),&j=\eta,\dots,d-1.\end{cases}$ (12)
This scale-aware MoE enhances representation capacity for low-scale features while maintaining computational efficiency for higher scales.
Localization-aware Score Calibration. Relying solely on classification scores can lead to low-quality bounding boxes being ranked higher than high-quality ones, resulting in suboptimal performance. This issue has been widely studied [18, 19, 20, 28, 12]. Unlike Cascade-DETR [28], which introduces a class-agnostic IoU score branch alongside the classification score, we propose a class-aware IoU score, inspired by VarifocalNet [18], to calibrate the classification score. During training, in contrast to Stable-DINO [19] and Align-DETR [12], the label assignment procedure excludes the IoU score, simplifying the prediction target and label assignment. We employ the VFL Loss [18] to learn the class-aware IoU score, where the target score is empirically lifted by $s_{target}=iou^{0.75}$. At inference, we compute a calibrated score by combining the classification and IoU scores:
$s_{\text{calib}}=s_{\text{cls}}^{\phi}\cdot s_{\text{iou}}^{1-\phi},$ (13)
where $s_{\text{cls}}\in[0,1]$ is the classification score, $s_{\text{pred}}\in[0,1]$ is the predicted IoU score, and $\phi$ is to balance the contributions of classification and localization confidence. This calibration is easy to use and does not affect the original model’s training process.
TABLE II: The performance on the COCO 2017 [90] validation set. All models are based on the ResNet-50 [91] backbone.
Model w/ Mr. DETR w/ Mr. DETR++ Epochs Queries AP AP50 AP75 APs APm APl
@@
Align-DETR [12] 24 900 52.4 (+1.1) 70.1 56.9 36.3 55.9 66.9
IV Experiments
IV-A Setup
IV-A1 Datasets and Evaluation
@@
We conduct extensive experiments on object detection, instance segmentation, and panoptic segmentation using three popular datasets (see App. -A): COCO 2017 [90], NuScenes [23], and Objects365 [22]. In line with existing research [11, 9, 8], we conduct evaluations on the COCO 2017 validation set, providing results in terms of standardized metrics, namely average precision (mAP, AP50, AP75) at various IoU thresholds. APs, APm, and APl refer to the average precision for small, medium, and large objects, respectively. For all experiments with 300 and 900 queries, we evaluate the performance according to top-100 and top-300 predictions, respectively. For the instance segmentation task, we report the mask AP and box AP, respectively. Panoptic segmentation is evaluated using the Panoptic Quality (PQ) metric [92]. We report the PQ for ‘thing’ and ‘stuff’ classes, respectively. Besides, we evaluate the box AP and mask AP for ‘thing’ classes.
IV-A2 Implement Details
@@
We conduct all experiments based on the ResNet-50 [91] and Swin-L [93] as the backbone pretrained on the ImageNet [94]. In all experiments, we apply the same data augmentations as [11, 9, 8, 13, 14]. AdamW [95] is employed as the optimizer, with the initial learning rate and weight decay established at 2e-4 and 1e-4, respectively. We utilize a batch size of 16 for training all the models. During training schedules of 12 and 24 epochs, the learning rate is reduced by a factor of 0.1 following the 11th and 20th epochs, respectively. In all experiments with instructive self-attention, we set the number of instruction tokens to 10 (see App. -B). We configure Mr. DETR++ with four experts, activating the top two experts for both the encoder and decoder (see App. -C). Within the one-to-many assigner, the hyper-parameters are configured as follows: $K=6$, $\alpha=0.3$, and $\tau=0.4$ (see App. -D). For models based on DINO [9], we use 100 contrastive denoising queries.
IV-B Object Detection
IV-B1 MS COCO 2017
We conduct extensive experiments on the COCO 2017 dataset. In Tab. II, we present the performance of our method across various baselines, including Deformable-DETR++ [11] with either 300 or 900 queries, DINO [9] and Align-DETR [12]. Specifically, using Deformable-DETR++ [11] with 300 queries, Mr. DETR achieves a 49.5% mAP, providing a 2.5% improvement without additional inference cost. It surpasses other variants, such as H-DETR [8], MS-DETR [14] and DAC-DETR [13], by approximately 0.7% in mAP. Mr. DETR++ reaches 51.0% mAP, outperforming the baseline model by 4%. Based on a strong baseline model like Deformable-DETR++ [11] with 900 queries, our Mr. DETR and Mr. DETR++ reach a 50.7% mAP and 51.8% under a 12-epoch training schedule, respectively. Extending to 24 epochs, Mr. DETR and Mr. DETR++ achieve a 51.4% and 52.0% mAP, outperforming the baseline by 1.6% and 2.2%, respectively. Compared to other variants with one-to-many auxiliary training, Mr. DETR++ exceeds DAC-DETR [13] and MS-DETR [14] by 1.5% and 1.1%, respectively.
TABLE III: The performance on the COCO 2017 [90] validation set. Our model is based on DINO [9] with IA-BCE loss [12]. All models are trained based on Swin-L [93] backbone for 12 epochs with 900 queries, except for H-DETR [8], which uses 300 queries.
Model AP AP50 AP75 APs APm APl
H-DETR [8] 55.9 75.2 61.0 39.1 59.9 72.2
DINO [9] 56.8 75.4 62.3 41.1 60.6 73.5
Co-DETR [57] 56.9 75.5 62.6 40.1 61.2 73.3
Salience-DETR [36] 56.5 75.0 61.5 40.2 61.2 72.8
Rank-DETR [20] 57.6 76.0 63.4 41.6 61.4 73.8
DAC-DETR [13] 57.3 75.7 62.7 40.1 61.5 74.4
EASE-DETR [29] 57.8 76.7 63.3 40.7 61.9 73.7
Stable-DINO [19] 57.7 75.7 63.4 39.8 62.0 74.7
Relation-DETR [30] 57.8 76.1 62.9 41.2 62.1 74.4
Mr. DETR (ours) 58.4 76.3 63.9 40.8 62.8 75.3
Mr. DETR++ (ours) 58.7 76.5 64.0 42.2 62.9 75.2
Based on DINO [9], our proposed Mr. DETR and Mr. DETR++ achieve 50.9% and 52.2% AP, respectively, over 12 epochs, surpassing variants such as Group-DETR [16], DAC-DETR [13] and MS-DETR [14]. When equipped with Mr. DETR and Mr. DETR++, Align-DETR yields AP improvements of 1.2% and 2.0%, respectively. Our experiments consistently demonstrate performance enhancements across different baseline models. Additional experiments utilizing the Swin-L [93] backbone, as shown in Tab. III, show that Mr. DETR and Mr. DETR++ attain 58.4% and 58.7% AP, respectively, outperforming other variants. These experimental results highlight the efficacy of our proposed method.
TABLE IV: Experiments on the large-scale Obejcts365 dataset [22]. Deformable-DETR++ (baseline) and our model use the ResNet-50 [91] backbone. All models are trained for 12 epochs with 900 queries.
Models AP AP50 AP75 APs APm APl
Baseline [11] 30.4 40.8 33.1 16.1 30.1 39.1
w/ Mr. DETR 32.7 (+2.3) 42.7 35.8 17.1 32.3 42.6
w/ Mr. DETR++ 34.9 (+4.5) 45.4 38.2 18.3 34.3 45.7
@@ IV-B1 MS COCO 2017
TABLE V: Experiments on the NuImages dataset [23]. Deformable-DETR++ (baseline) and our models use the ResNet-50 [91] backbone. All models are trained for 12 epochs with 900 queries.
Models AP AP50 AP75 APs APm APl
Baseline [11] 48.5 75.1 52.3 27.4 46.1 63.0
w/ Mr. DETR 51.2 (+2.7) 76.4 55.8 28.3 48.1 66.4
w/ Mr. DETR++ 52.3 (+3.8) 75.8 56.5 25.5 49.0 69.0
TABLE VI: Instance segmentation results on the COCO 2017 [90] validation set. All experiments are based on Deformable-DETR++ [11] with 300 queries and ResNet-50 [91] as backbone.
Models Epochs Mask AP Box AP
Deformable-DETR++ [11] 12 32.4 46.5
w/ Mr. DETR 12 36.0 (+3.6) 49.5 (+3.0)
w/ Mr. DETR++ 12 37.7 (+5.3) 50.8 (+4.3)
Deformable-DETR++ [11] 24 35.1 48.6
w/ Mr. DETR 24 37.6 (+2.5) 50.3 (+1.7)
w/ Mr. DETR++ 24 39.0 (+3.9) 51.1 (+2.5)
IV-B2 Objects365
To further verify the scalability of our method on the large-scale dataset, we conduct experiments on the Objects365 [22] dataset using Deformable-DETR++ [11] model with 900 queries. We train the baseline model and our method for 4 epochs only. The initial learning rate is set to 2e-4 and decays at the third epoch. Other training settings are the same as the model trained on the COCO 2017 dataset [90]. As reported in Tab. IV, Mr. DETR and Mr. DETR++ achieve 2.3% and 4.5% improvement over the baseline model in terms of AP, corroborating the effectiveness of our method.
IV-B3 NuImages
We evaluate our approach on the NuImages [23] dataset, which encompasses diverse traffic scenarios. The experimental results are presented in Tab. V. Mr. DETR achieves an AP of 51.2%, surpassing the baseline model by 2.7%. Mr. DETR++ further improves performance, outperforming the baseline by 3.8% AP, highlighting the effectiveness of our method. These results on the NuImages dataset underscore the potential of our approach for autonomous driving applications.
(a) Evaluation results of box
(b) Evaluation results of mask
IV-C Instance Segmentation